# ON THE AVERAGE ULTRAVIOLET EMISSION LINE SPECTRA OF HIGH-REDSHIFT GALAXIES: HOT AND COLD, CARBON-POOR, NITROGEN-MODEST, AND OOZING IONIZING PHOTONS

### v1

@@ Abstract
We determine the spectroscopic properties of ostensibly star-forming galaxies at redshifts () using prism spectroscopy from JWST/NIRSpec. With rest-wavelength coverage between Ly and [S ii] in the optical, we stack spectra as a function of nebular conditions, and compare UV spectral properties with stellar age. This reveals UV lines of N iii], N iv], C iii], C iv, He ii, and O iii] in the average high- galaxy. All UV lines are more intense in younger starbursts. We measure electron temperatures from the collisionally excited [O iii] line ratios, finding K for the O++ regions. We also detect a significant nebular Balmer Jump from which we estimate only K. Accounting for typical temperature offsets between zones bearing doubly and singly ionized oxygen, these two temperatures remain discrepant by around 40%. We use the [O iii] temperatures to estimate abundances of carbon, nitrogen, and oxygen. We find that log(C/O) is consistently , with no evolution of C/O with metallicity or stellar age. The average spectra are mildly enhanced in Nitrogen, with higher N/O than low- starbursts, but are less enhanced than samples of high- galaxies with visible UV N iii] and N iv]. Whatever processes produce the N-enhancement in the individual galaxies must also be ongoing, at lower levels, in the median galaxy in the early Universe. The strongest starbursts are a source of significant ionizing emission: ionizing photon production efficiencies reach Hz erg-1, and show multiple signatures of high Lyman continuum escape, including Mg ii escape fractions nearing 100%, significant deficits in [S ii] emission, high degrees of ionization, and blue UV colors.
@@ 1 Introduction
The James Webb Space Telescope (JWST) has revolutionized our understanding of early galaxy formation. Over the last two years, data from the telescope have updated our view on the number density of galaxies at the dawn of cosmic star-formation (e.g. Bouwens et al., 2023; Donnan et al., 2023; Finkelstein et al., 2024; Harikane et al., 2023a; Robertson et al., 2024), and enabled the intriguing discovery of massive-black-hole galaxy hosts (e.g. Maiolino et al., 2023; Kocevski et al., 2024; Kokorev et al., 2024; Scholtz et al., 2024), and dusty, quenched systems at high- (Barrufet et al., 2024; Carnall et al., 2024; Weibel et al., 2024). It has allowed us to witness the early chemical enrichment of galaxies (e.g. Arellano-Córdova et al., 2022; Curti et al., 2023; Schaerer et al., 2022a; Trump et al., 2023; Stiavelli et al., 2023), as well as the ionization history and topology of the Universe using the Ly emission line (e.g. Hayes & Scarlata, 2023; Bruton et al., 2023a, b; Tang et al., 2023, 2024; Witstok et al., 2024; Napolitano et al., 2024).
With the advent of NIRSpec spectroscopy (Jakobsen et al., 2022), the first statistical samples of high- galaxies with rest-optical spectroscopy have been assembled. Emission line ratio diagrams (e.g. Backhaus et al., 2024; Cameron et al., 2023; Mascia et al., 2023; Shapley et al., 2024; Sanders et al., 2024), suggest gas to be highly excited, due to illumination by a hard ionizing spectrum. Moreover, access to emission lines whose ratios are temperature-sensitive has enabled the determination of robust estimates of oxygen abundances, facilitating new metallicity calibrations (e.g. Sanders et al., 2024; Curti et al., 2024) and therefore the determination of the mass-metallicity relation (MZR) for the first time at (e.g. He et al., 2024; Nakajima et al., 2023; Chemerynska et al., 2024). At these low masses (or high redshifts), not only does the normalization of the MZR decrease, but its slope also flattens, hinting a change in the wind mechanisms from energy (high-mass) to momentum-driven outflows (in the dwarf regime), with substantial implications for our understanding of the baryon cycle of early galaxies.
@@
JWST has made substantial progress in constraining the chemical enrichment of high- galaxies. Carbon abundances (usually given as C/O) have been studied from the ground (e.g. Erb et al., 2010; Stark et al., 2014; Mainali et al., 2020; Matthee et al., 2021; Citro et al., 2024) upon which JWST could build (e.g. Arellano-Córdova et al., 2022; Curti et al., 2023; Isobe et al., 2023; Jones et al., 2023), usually finding C/O ratios consistent with the local universe at fixed (and usually low) oxygen abundance. Perhaps even more interestingly, JWST spectra have identified a handful of ‘nitrogen-loud’ galaxies, of which seven are currently known: GNz11 (Bunker et al., 2023), GLASS_150008 (Isobe et al., 2023), CEERS-1019 (Marques-Chaves et al., 2024), GS3073 (Ji et al., 2024, originally found by Vanzella et al. 2010), GHZ2 (Castellano et al., 2024), RXCJ2248-4431 (Topping et al., 2024), and GN-z9p4 (Schaerer et al., 2024). These objects are peculiar, and seem to be absent at lower redshifts, although the noteworthy exception of Mrk 996 implies this is not uniquely true (James et al., 2009). While nitrogen should follow qualitatively similar primary (core-collapse supernova) and secondary (low-intermediate mass late stages/AGB) enrichment patterns as carbon, it is difficult to explain the combined finding that carbon is lacking but nitrogen is enhanced (e.g. Berg et al., 2019b, 2020). These objects have therefore triggered a lot of speculation about the reasons for their nitrogen luminosity, including exotic scenarios such as supermassive stars produced by stellar mergers in dense star clusters (e.g. Charbonnel et al., 2023), AGN activity (e.g. Maiolino et al., 2024), redshift 11 Tidal Disruption Events (Cameron et al., 2023) or carefully balanced recent star-formation histories (Kobayashi & Ferrara, 2024).
@@
Finally, many of the objects discussed above for which detailed emission line spectroscopy has been obtained lie at redshifts where the Universe is demonstrated to be substantially neutral. Some of the same features used for classification and abundance inferences are also key estimators of both the intrinsic ionizing power of a galaxy (e.g. the Balmer series of emission lines) and the escape of ionizing radiation (e.g. [O iii]5008, [O ii]3727, [S ii]6717,31, Mg ii2800). At these redshifts, the Universe is fully opaque to ionizing radiation (LyC, Inoue et al., 2014) and the ioniziong power of galaxies can never be observed directly. Thus the mentioned emission lines are critical observables in the reionization epoch, as they provide indirect diagnostics of the amount of ionizing radiation escaping galaxies (, E.g. Leitet et al., 2013; Izotov et al., 2016; Flury et al., 2022a; Chisholm et al., 2022).
@@
As exciting as these results are, they are mostly restricted to small samples of individual galaxies (usually one per paper, where nitrogen enhancement is concerned) that are identified by carefully inspecting archival spectra and publishing the gems. Given the smaller rest-frame EWs of these UV emission lines (compared to their optical counterparts) they are harder to detect – this causes JWST to lag somewhat in the delivery of UV-based inferences compared to more traditional methods using optical lines. The uncertainties on these metal abundances are often large, and it is likely that there is little to be gained by publishing large numbers of limits on (C/O) and (N/O) for the plethora of non-detections. Thus we have little insight into the distribution of these abundances across the majority of the iceberg that remains submerged.
@@
In this paper we adopt a similar approach, and perform a stacking analysis of galaxies at for which NIRSpec prism observations have been obtained. We use the DAWN JWST Archive (DJA111https://dawn-cph.github.io/dja ; see Heintz et al., 2024) to obtain a large sample of galaxies at , which we then combine in subsets of galaxies per bin, arriving at total effective integration times of several million seconds per spectral pixel. This extremely deep spectroscopy reveals the UV line emission from the galaxies, which we then measure to explore astrophysical processes in concert with the optical lines. We study emission lines of [C ii], C iii], C iv, He ii, O iii], N iii] and N iv] in the UV, as well as many strong lines in the optical range. We make a concerted effort to understand the nebular abundances, the variation in their patterns, and evolution with the youth of the star formation episode. In Section 2 we present the sample selection, data, and data analysis methods. In Section 3 we show the average spectra of high- galaxies, describe the main trends with star-formation evolutionary state, and quantify the lines ratios and EWs. In Section 4 we calculate electron temperatures using both collisional lines of [O iii] and the nebular Balmer Jump, and estimate the abundances of oxygen, carbon, and nitrogen. We describe measurements related to the production and escape of ionizing radiation ( and ) in Section 5. In Section 6 we discuss how these observations relate to other efforts to understand the buildup and chemical enrichment of galaxies, and in Section 7 we present our concluding remarks. Throughout we assume a cosmology of , and the AB magnitude system (Oke & Gunn, 1983).
@@ 2.1 Sample Selection
The sample is drawn from the DJA archive, similar to Heintz et al. (2024, see also and ). We require a large number of galaxies and complete spectral coverage between Ly and H, which is satisfied by the samples for which PRISM/CLEAR spectroscopy have been obtained. The prism has a wavelength range of 0.6–5.3 m, with a resolving power of , depending upon the wavelength. Our main aim is to obtain the deepest possible spectrum of the UV nebular lines, so we set a lower limit of , which shifts the Ly break and emission line into the wavelength coverage of NIRSpec/PRISM. Wishing to study galaxies to the largest possible distance, we set no upper limit in redshift. This yields a sample of 1115 galaxies. They are drawn from the multitude of JWST spectroscopic campaigns from all cycles, including General Observing campaings, Early Release Science, Guaranteed Time Observations, and Director’s Discretionary Time, and combine all kinds of different selection techniques.
@@ 2.2 Continuum Modeling and Stacking
Our goal is to study line emission, so the stellar continuum must either be removed by subtraction or normalized. We make use of both procedures in this paper: subtraction of continuum conserves line fluxes, allowing them to be measured and line ratios calculated; normalization of the continuum conserves equivalent width. The best approach would be to model the stellar continuum with population synthesis models (e.g Hayes et al., 2023), but these individual spectra have neither the resolution nor signal-to-noise ratio (S/N) to make such fitting meaningful. We therefore adopt simple polynomial representations of the continuum. Such continuous functions will not capture spectral discontinuities such as the Balmer/4000Å break, so we perform this functional fitting over two separate wavelength regions for each object: the UV (at Å to avoid the Balmer break) and optical (at Å to avoid the [O ii] emission line).
@@
To improve the redshift precision for stacking, we make a first round of fits to emission lines in the individual spectra. We fit single Gaussian functions to the optical emission lines of [O ii], H, [O iii], and H in the continuum-subtracted spectra. The [O iii]4960,5008 doublet is represented by a pair of emission lines with constrained wavelength and flux ratios. Not all the lines are fit at all redshifts, since H is lost at and [O iii] is lost at . We deem it unreliable to base redshifts purely upon [O ii] line, so limit our study to galaxies where [O iii], H, and [O ii] are available, giving us a de facto cap of . We deem a fit successful if we obtain measurements in any optical emission line, which is the effective selection function of our study within the DJA archive. Removing a handful of galaxies for which this line fitting fails, we are left with a sample of 1035 galaxies. We show the basic properties of this sample in Figure 1 and provide more details follow in the coming sections.
With redshifts based upon spectral lines, we shift each spectrum into the restframe, and resample them onto a common rest wavelength grid spaced at 5Å. We transform flux density to luminosity density , by multiplying by , where is the luminosity distance. We produce four grids of the restframe spectra: (1.) those with no normalization so we can investigate the properties of the spectra in absolute luminosities; (2.) those normalized by the monochromatic continuum luminosity at 3500Å, so we can investigate changes in colors; (3.) those with the local continuum subtracted at each wavelength (using the fitted polynomial), so we can measure emission lines luminosities in their absolute units; and (4.) those with the variable continuum normalized (again at each wavelength using the same polynomial fit), so that equivalent width can be derived. We compute uncertainties on the stacked spectra using a bootstrap-with-replacement simulation with 1000 realizations.
@@
In this procedure we have recorded a number of quantities. We store (1500Å) and (over the 1500–2200Å range) from the power-law fitting to the UV portion of the spectrum, and the fluxes and EWs in [O ii], H, [O iii], and H. We use these tabulated variables as vectors along which we sub-divide the sample into bins, and stack each of the grids of spectra. Thus we can investigate changes in the average spectra as a function of these properties. In the main part of this paper we use the optical [O iii] equivalent width () as our independent variable, because it is the strongest spectral feature that we have in the optical spectrum, and closely traces the youth of the ongoing star formation episode as it is primarily dependent upon the number of ionizing photons produced with energies eV. We divide the galaxies into five approximately equally spaced bins, at boundaries of 200, 500, 1000, 1500 Å, with total integration times in each stack reaching several million seconds. The basic properties of these individual stacks are shown in Table 1.

### v2

# ON THE AVERAGE ULTRAVIOLET EMISSION LINE SPECTRA OF HIGH-REDSHIFT GALAXIES: HOT AND COLD, CARBON-POOR, NITROGEN MODEST, AND OOZING IONIZING PHOTONS

@@ Abstract
We determine the spectroscopic properties of ostensibly star-forming galaxies at redshifts () using prism spectroscopy from JWST/NIRSpec. With rest-wavelength coverage between Ly and [S ii] in the optical, we stack spectra as a function of nebular conditions, and compare UV spectral properties with stellar age. This reveals UV lines of N iii], N iv], C iii], C iv, He ii, and O iii] in the average high- galaxy. All the UV lines are more intense in younger starbursts. We measure electron temperatures from the collisionally excited [O iii] line ratios, finding K for the O++ regions. We also detect a significant nebular Balmer Jump, from which we estimate only K. Accounting for typical temperature offsets between zones bearing doubly and singly ionized oxygen, these two temperatures remain discrepant by around 40%. We use the [O iii] temperatures to estimate abundances of carbon, nitrogen, and oxygen. We find that log(C/O) is consistently , with no evolution of C/O with metallicity or stellar age. The average spectra are mildly enhanced in Nitrogen, with higher N/O than low- starbursts, but are less enhanced than samples of recently reported, high-, extreme galaxies that show N iii] and N iv] emission in the UV. Whatever processes produce the N-enhancement in the individual galaxies must also be ongoing, at lower levels, in the median galaxy in the early Universe. The strongest starbursts are a source of significant ionizing emission: ionizing photon production efficiencies reach Hz erg-1, and show multiple signatures of high Lyman continuum escape, including Mg ii escape fractions nearing 100%, significant deficits in [S ii] emission, high degrees of ionization, and blue UV colors.
@@ 1 Introduction
The James Webb Space Telescope (JWST) has revolutionized our understanding of early galaxy formation. Over the last two years, data from the telescope have updated our view on the number density of galaxies at the dawn of cosmic star-formation (e.g. Bouwens et al., 2023; Donnan et al., 2023; Finkelstein et al., 2024; Harikane et al., 2023a; Robertson et al., 2024), and enabled the intriguing discovery of massive-black-hole galaxy hosts (e.g. Maiolino et al., 2023; Kocevski et al., 2024; Kokorev et al., 2024; Scholtz et al., 2024), as well as dusty, quenched systems (Barrufet et al., 2024; Carnall et al., 2024; Weibel et al., 2024). It has allowed us to witness the early chemical enrichment of galaxies (e.g. Arellano-Córdova et al., 2022; Curti et al., 2023; Schaerer et al., 2022a; Trump et al., 2023; Stiavelli et al., 2023), as well as the ionization history and topology of the Universe using the Ly emission line (e.g. Hayes & Scarlata, 2023; Bruton et al., 2023a, b; Tang et al., 2023, 2024; Witstok et al., 2024; Napolitano et al., 2024a).
With the advent of NIRSpec spectroscopy (Jakobsen et al., 2022), the first statistical samples of high- galaxies with rest-optical spectroscopy have been assembled. Emission line ratio diagrams (e.g. Backhaus et al., 2024; Cameron et al., 2023; Mascia et al., 2023; Shapley et al., 2024; Sanders et al., 2024), suggest gas to be highly excited, due to illumination by a hard ionizing spectrum. Moreover, access to emission lines whose ratios are temperature-sensitive has enabled the determination of robust estimates of oxygen abundances, facilitating new metallicity calibrations (e.g. Sanders et al., 2024; Curti et al., 2024) and therefore the determination of the mass-metallicity relation (MZR) for the first time at (e.g. He et al., 2024; Nakajima et al., 2023; Chemerynska et al., 2024). At these low masses (or high redshifts), not only does the normalization of the MZR decrease, but its slope also flattens, which may hold substantial implications for our understanding of the baryon cycle of early galaxies.
@@
JWST has made substantial progress in constraining the chemical enrichment of high- galaxies. Carbon abundances (usually given as C/O) have been studied from the ground (e.g. Erb et al., 2010; Stark et al., 2014; Mainali et al., 2020; Matthee et al., 2021; Citro et al., 2024) upon which JWST could build (e.g. Arellano-Córdova et al., 2022; Curti et al., 2023; Isobe et al., 2023; Jones et al., 2023), usually finding C/O ratios consistent with the local universe at fixed (and usually low) oxygen abundance. Perhaps even more interestingly, JWST spectra have identified a handful of ‘nitrogen-loud’ galaxies, of which seven are currently known: GNz11 (Bunker et al., 2023), GLASS_150008 (Isobe et al., 2023), CEERS-1019 (Marques-Chaves et al., 2024), GS3073 (Ji et al., 2024, originally found by Vanzella et al. 2010), GHZ2 (Castellano et al., 2024), GHZ9 (Napolitano et al., 2024b), RXCJ2248-4431 (Topping et al., 2024), and GN-z9p4 (Schaerer et al., 2024). These objects are peculiar, and seem to be absent at lower redshifts, although the noteworthy exception of Mrk 996 (James et al., 2009) implies this is not uniquely true. While nitrogen should follow qualitatively similar primary (core-collapse supernova) and secondary (low-intermediate mass late stages/AGB) enrichment patterns as carbon, it is difficult to explain the combined finding that carbon is lacking but nitrogen is enhanced (e.g. Berg et al., 2019b, 2020). These objects have therefore triggered a lot of speculation about the reasons for their nitrogen luminosity, including exotic scenarios such as supermassive stars produced by stellar mergers in dense star clusters (e.g. Charbonnel et al., 2023), AGN activity (e.g. Maiolino et al., 2024), Tidal Disruption Events (Cameron et al., 2023) or carefully balanced recent star-formation histories (Kobayashi & Ferrara, 2024).
@@
Finally, many of the objects discussed above for which detailed emission line spectroscopy has been obtained lie at redshifts where the Universe is demonstrated to be substantially neutral. Some of the same features used for classification and abundance inferences are also key estimators of both the intrinsic ionizing power of a galaxy (e.g. the Balmer series of emission lines) and the escape of ionizing radiation (e.g. [O iii]5008, [O ii]3727, [S ii]6717,31, Mg ii2800). At these redshifts, the Universe is fully opaque to ionizing radiation (LyC, Inoue et al., 2014) and the ionizing power of galaxies can never be observed directly. Thus the mentioned emission lines are critical observables in the reionization epoch, as they provide indirect diagnostics of the amount of ionizing radiation escaping galaxies (, E.g. Leitet et al., 2013; Izotov et al., 2016; Flury et al., 2022a; Chisholm et al., 2022).
@@
As exciting as these results are, they are mostly restricted to small samples of individual galaxies (usually one per paper, where nitrogen enhancement is concerned) that are identified by carefully inspecting archival spectra and publishing the gems. Given the small rest-frame EWs of these UV emission lines (compared to their optical counterparts) they are harder to detect – this causes JWST to lag somewhat in the delivery of UV-based inferences compared to more traditional methods using stronger lines in the optical. The uncertainties on these metal abundances are often large, and it is likely that there is little to be gained by publishing large numbers of limits on (C/O) and (N/O) for the plethora of non-detections. Thus we have little insight into the distribution of these abundances across the majority of the iceberg that remains submerged.
@@
In this paper we adopt a similar approach, and perform a stacking analysis of galaxies at for which NIRSpec prism observations have been obtained. We use the DAWN JWST Archive (DJA111https://dawn-cph.github.io/dja ; see Heintz et al., 2024) to obtain a large sample of galaxy spectra at , which we then combine in subsets of galaxies per bin, arriving at total effective integration times of several million seconds per spectral pixel. This extremely deep spectroscopy reveals the UV line emission from the galaxies, which we then measure to explore astrophysical processes in concert with the optical lines. We study emission lines of [C ii], C iii], C iv, He ii, O iii], N iii] and N iv] in the UV, as well as many strong lines in the optical range. We make a concerted effort to understand: (1.) the nebular gas temperatures; (2.) the associated nebular abundances of C, N, and O and variation in their patterns; (3.) the production efficiency of ionizing radiation; and (4.) the ionizing escape fraction. With information on the stellar population age, we investigate all this as a function of stellar evolutionary stage.
The paper proceeds as follows. In Section 2 we present the sample selection (§2.1), data processing and spectral coaddition (§2.2), and emission line measurement methods (§2.3). In Section 3 we show the average spectra of high- galaxies, describe the main trends with star-formation evolutionary state, quantify the lines ratios and EWs, and contrast the stacked high- results with well-studied low- samples – we do this qualitatively in §3.1 and quantitatively in §3.2. In Section 4 we calculate electron temperatures using both the Balmer Jump (§4.1.1) and collisional lines of [O iii] (§4.1.2), estimate the abundances of oxygen, carbon, and nitrogen (§4.2) and discuss the impact of our assumptions on these results (§4.3). We describe measurements related to the production and escape of ionizing radiation ( and ) in Section 5. In Section 6 we discuss how these observations inform our view of the ionizing conditions in early galaxies (§6.1), gas temperatures (§6.2), existing efforts to understand the buildup and chemical enrichment of galaxies (§6.3), and the emission of ionizing radiation (§6.4). In Section 7 we present our concluding remarks. Throughout we assume a cosmology of , and the AB magnitude system (Oke & Gunn, 1983).
@@ 2.1 Sample Selection
The sample is drawn from the DJA archive, similar to Heintz et al. (2024, see also and ). We require a large number of galaxies and complete spectral coverage between Ly and H, which is satisfied by the samples for which PRISM/CLEAR spectroscopy have been obtained. The prism has a wavelength range of 0.6–5.3 m, with a resolving power of , depending upon the wavelength. Our main aim is to obtain the deepest possible spectrum of the UV nebular lines, so we set a lower limit of , which shifts the Ly break and emission line into the wavelength coverage of NIRSpec/PRISM. Wishing to study galaxies to the largest possible distance, we set no upper limit in redshift. Each object in the DJA spectroscopic archive includes a grade flag, which identifies the quality of the spectrum; we take only spectra with grade , corresponding to the highest quality sources, although all spectra were also visually inspected. This yields a sample of 1115 galaxies. They are drawn from the multitude of JWST spectroscopic campaigns from all cycles, including General Observing campaigns, Early Release Science, Guaranteed Time Observations, and Director’s Discretionary Time, and combine all kinds of different selection techniques.
@@ 2.2 Continuum Modeling and Stacking
Our goal is to study line emission, so the stellar continuum must either be removed by subtraction or normalized. We make use of both procedures in this paper: subtraction of continuum conserves line fluxes, allowing them to be measured and line ratios calculated; normalization of the continuum conserves equivalent width of each line. The best approach would be to model the stellar continuum with population synthesis models (e.g Hayes et al., 2023), but these individual spectra have neither the resolution nor signal-to-noise ratio (S/N) to make such fitting meaningful. We therefore adopt simple polynomial representations of the continuum. Such continuous functions will not capture spectral discontinuities such as the Balmer/4000Å break, so we perform this functional fitting over two separate wavelength regions for each object: the UV (at Å to avoid the Balmer break) and optical (at Å to avoid the [O ii] emission line).
@@
To verify, and in some cases improve, the redshift precision for the stacking we make a first round of fits to emission lines in each spectrum. We assume that most of the important information for redshifts will be present in the strong optical emission lines of [O ii], H, [O iii], and H, and that most of these lines could be visible for individual galaxies. We perform a simultaneous fit to these emission lines using the same engine as used in Hayes et al. (2023). Fixing the redshifts of all features to be the same, we fit single Gaussian functions for [O ii], H, and H (ignoring possible contamination from the adjacent [N ii] doublet), and a double Gaussian to the [O iii] feature in which the 4960Å emission line is fixed to be 1/3 the strength of the 5008Å line. The amplitude for each line is its only individual free parameter and, thus, the four features are fit to provide a single redshift. We note that the H line is lost from the red end of NIRSpec at and [O iii]+H are lost at . We deem it unreliable to base redshifts purely upon [O ii] line, so limit our study to galaxies where [O iii], H, and [O ii] are available, giving us a de facto cap of . We declare a fit to be ‘successful’ if we obtain measurements in any of the features, which provides the effective selection function of our study within the DJA archive. All fits were inspected by eye. Removing a handful of galaxies for which this line fitting fails, we are left with a sample of 1035 galaxies. We show the basic properties of this sample in Figure 1 and provide more details in the coming sections.
With redshifts based upon spectral lines, we shift each spectrum into the restframe, and resample them onto a common rest wavelength grid spaced at 5Å. We transform flux density to luminosity density , by multiplying by , where is the luminosity distance. We produce four grids of the restframe spectra: (1.) those with no normalization so we can investigate the properties of the spectra in absolute luminosities; (2.) those normalized by the monochromatic continuum luminosity at 3500Å, so we can investigate changes in continuum colors; (3.) those with the local continuum subtracted at each wavelength (using the fitted polynomial), so we can measure emission lines luminosities in their absolute units; and (4.) those with the variable continuum normalized (again at each wavelength using the same polynomial fit), so that equivalent width can be derived. We compute uncertainties on the stacked spectra using a bootstrap-with-replacement simulation with 1000 realizations.
@@
In this procedure we have recorded a number of quantities. We store (1500Å) and (over the 1500–2200Å range) from the power-law fitting to the UV portion of the spectrum, and the fluxes and EWs in [O ii], H, [O iii], and H. We use these tabulated variables as vectors along which we sub-divide the sample into bins, and stack each of the grids of spectra. Thus we can investigate changes in the average spectra as a function of these properties. In this paper we use the optical [O iii] equivalent width () as our independent variable, because it is the strongest spectral feature that we have in the optical spectrum. closely traces the youth of the ongoing star formation episode as at most metallicities it is primarily dependent upon the number of ionizing photons produced per unit time333This relationship must break at very low metallicities, where the [O iii] emission becomes sub-dominant to H, as has been potentially identified in very faint galaxies by Endsley et al. (2024), although see also Begley et al. (2024). In the Endsley et al. (2024) sample, this effect becomes most prominent in their , of which there are very few such galaxies in our sample (Figure 1); while this must affect our work to some degree, we expect the impact to be minor (especially given the median stacking performed).. We divide the galaxies into five approximately equally spaced bins, at boundaries of 200, 500, 1000, 1500 Å, with total integration times in each stack reaching several million seconds. The basic properties of these individual stacks are shown in Table 1.